OpenStudy (anonymous):

A hydrogen atom has an electron of mass m and a proton of mass M, M>>m. The reduced mass is defined to be: u= mM/(m+M) Given u as a series in m/M: The 1st order correction term is defined by including the linear term by no higher order terms. If m = M/1836, by what percentage does including the 1st order correction change the estimate u=m?

OpenStudy (perl):

did you do the series for mM / (m + M )

OpenStudy (perl):

u= m * M / (m+M) u = m* M * ( m + M ) ^ -1 u = m*M * [ M ( m/M + 1) ] ^ -1 u = m * M * M^-1 ( m/M + 1 )^-1 u = m * ( 1 + m/M) ^-1

OpenStudy (perl):

now we can use the taylor series ( 1 + x) ^-1 = 1 - x + x^2 - x^3 + ...

OpenStudy (perl):

u = m * ( 1 + m / M ) ^-1 u = m * [ 1 - m/M + (m/M)^2 - (m/M)^3 + ...

OpenStudy (perl):

now the directions say , use linear correction (no higher orders )

OpenStudy (perl):

use 1st order correction*

OpenStudy (perl):

u = m * [ 1 - m/M + (m/M)^2 - (m/M)^3 + ... 1st order correction u ~ m * ( 1 - m/M)

OpenStudy (perl):

is it possible you can take a screen shot of the question, sometimes seems odd

OpenStudy (perl):

is that the letter mu, not u

OpenStudy (anonymous):

Yes, is is mu, not u

OpenStudy (anonymous):

this was a problem I got in class, I can't take a picture of it